If b= 16.3 ft. and a=21.7 ft., find /_A. Thanks so much for your help.
Answer by stanbon(75887)

If b= 16.3 ft. and a=21.7 ft., find /_A.
---------
Draw a right triangle.
Label the hypotenuse "c"; the side opposite angle A as "a"; the remaining
side "b".
----------
EQUATION:
tan(A) = a/b
tan(A) = 21.7/16.3 = 1.33129
angle A = tan^-1(1.33129) = 53.088 degrees
==============
Cheers,
Stan H.